hi,
so, my question to you is that :
what will be the answer to this question?
how many spheres will fit in a 30cm*20cm*10cm tin?
radius of the sphere is: 4.88m
so please help me with this question... thank you soooooo much! and please also explain how you got the answer! thank you! :)

Answer by mananth(16949)

how many spheres will fit in a 30cm*20cm*10cm tin?
radius of the sphere is: 4.88m

volume of tin / volume of each sphere = number of spheres
volume of tin = 30*20*10= 6000 cm^3
I assume the radius is 4.88cm
volume of sphere = (4/3) * pi*4.88^3
=486.55
6000/486.55
= 12 spheres